OpenStudy (anonymous):

Finding real or imaginary solutions of equations by factoring need help!!! 64x^3=-8

OpenStudy (anonymous):

64^3=262144

OpenStudy (anonymous):

262144 does not = -8

OpenStudy (anonymous):

please medal

OpenStudy (anonymous):

That isn't right... there is some kind of technique that I don't understand and I need someone to help me

OpenStudy (anonymous):

oh opps i forgot a step

OpenStudy (anonymous):

once you have 262144x= -8, divide both sides by -8

OpenStudy (anonymous):

x= - 32768

OpenStudy (anonymous):

No.. that's not the kind of algebra I'm working on... I know how to do that. This lesson is something with polynomials and a bunch of crap that doesn't make sense.

OpenStudy (solomonzelman):

I don't see why the solutions should be imaginary. They are not!

OpenStudy (solomonzelman):

\(\large\color{blue}{ 64x^3=-8 }\) \(\large\color{blue}{ 4^3x^3=-8 }\) \(\large\color{blue}{ (4x)^3=-8 }\) \(\large\color{blue}{ (4x)^3=(-2)^3 }\) \(\large\color{blue}{ 4x=-2 }\) \(\large\color{red}{ x=-\frac{1}{2} }\)